OpenStudy (anonymous):

I NEED HELP IN FACTORIZING THE SUM OF TWO DIFFERENT CUBES ....... X^3+8=X^3+2^3=(X+2)(X^2-X times +2^2

OpenStudy (angela210793):

X^3+8=(x+2)*(x^2-2x+4) x=-2 is the root

OpenStudy (anonymous):

thnxxxxxxxxxx angela

OpenStudy (angela210793):

uw :)